# A new parametrization of Hubble function and Hubble tension

### v2

@@ IV Results
In this section, We discuss the results in Sec III. We leveraged multiple observational datasets, including Hubble, Hubble+Pantheon samples and Hubble+Pantheon+BAO data [81]. Employing the MCMC method, we finally constrained the model parameters $H_{0}$, $\Omega_{\rm{m0}}$, and $\alpha$.
Firstly, the Hubble dataset, consisting of 62 data points, was utilized to quantify the historical expansion of the universe. In Figure 1, the best-fitting values for the model parameters were determined as $H_{0}=65.4^{+2.0}_{-2.2}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.266\pm{0.013}$ , and $\alpha=0.71\pm{0.31}$. It can be observed that the fitting parameters $H_{0}$ have relatively large uncertainties, and a significant contributing factor is the limited amount of data. In Figure 2, we can find that our model is relatively close to the $\Lambda$CDM model, and it fits well with the observational data. Next, in order to reduce the uncertainties in the parameters, we have incorporated 1701 observational data points from the Pantheon dataset [71]. Figure 3 presents the fitting results: $H_{0}=67.2\pm{1.7}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}$ = $0.264\pm{0.012}$ and $\alpha$ =$0.47\pm{0.26}$. It is observed that the error in $H_{0}$ has decreased. In Figure 4, $\alpha$CDM is seen to fit well with the observational data from Pantheon and is close to the $\Lambda$CDM model. Next, we included an additional 6 BAO data points and obtained the fitting results: $H_{0}=67.5^{+1.3}_{-1.6}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.2764\pm{0.0094}$, and $\alpha=0.33\pm{0.22}$. In Table 5, we can find with the increase in the observational data size, the constraints on the parameters $H_{0}$, $\Omega_{\rm{m0}}$, and $\alpha$ become progressively more accurate. To validate the model’s efficacy, we compared error bar plots for Hubble and Pantheon datasets with the $\Lambda$CDM model. Consistently, the model fits the observed dataset very well.
@@ V Conclusions
In conclusion, we proposed a new Hubble parameterization method and constrained the model parameters using observations from Hubble, Pantheon, and BAO. This model is validated with Hubble, Pantheon, and BAO data, providing best-fit values for $H_{0}=67.5^{+1.3}_{-1.6}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.2764\pm{0.0094}$, and $\alpha=0.33\pm{0.22}$, consistent well with the Planck 2018 data ($H_{0}=67.4\pm 0.5$ Km/s/Mpc [2]), but deviating from Cepheid-supernova observation ($H_{0}=73.04\pm 1.04$ $\text{km s}^{-1}\text{Mpc}^{-1}$ [7]) by more than 4 $\sigma$. The Hubble tension was mainly triggered with the higher Hubble constant ($H_{0}=74.0\pm 1.04$ $\text{km s}^{-1}\text{Mpc}^{-1}$ [7]) estimated from the local Cepheid–type Ia supernova distance ladder being at odds with the lower value extrapolated from CMB data, assuming the standard $\Lambda$CDM cosmological model ($H_{0}=67.4\pm 0.5$ $\text{km s}^{-1}\text{Mpc}^{-1}$ [2]). However, our analyses indicate that the Hubble tension may also exist between the Hubble+Pantheon+BAO data and the measurements from local Cepheid–type Ia supernova distance ladder if taking the parametrization (5). We also constrained $w$CDM with the same datasets for comparison and found that $\alpha$CDM is consistent with $w$CDM within 1 $\sigma$ error, indicating that the impact of radiation is relatively weak.

### v3

@@ IV Results
In this section, We discuss the results in Sec III. We leveraged multiple observational datasets, including Hubble, Hubble+Pantheon samples and Hubble+Pantheon+BAO data Giostri et al. (2012). Employing the MCMC method, we finally constrained the model parameters $H_{0}$, $\Omega_{\rm{m0}}$, and $\alpha$.
Firstly, the Hubble dataset, consisting of 62 data points, was utilized to quantify the historical expansion of the universe. In Figure 1, the best-fitting values for the model parameters were determined as $H_{0}=65.4^{+2.0}_{-2.2}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.266\pm{0.013}$ , and $\alpha=0.71\pm{0.31}$. It can be observed that the fitting parameters $H_{0}$ have relatively large uncertainties, and a significant contributing factor is the limited amount of data. In Figure 2, we can find that our model is relatively close to the $\Lambda$CDM model, and it fits well with the observational data. Next, in order to reduce the uncertainties in the parameters, we have incorporated 1701 observational data points from the Pantheon dataset Perivolaropoulos and Skara (2022). Figure 3 presents the fitting results: $H_{0}=67.2\pm{1.7}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}$ = $0.264\pm{0.012}$ and $\alpha$ =$0.47\pm{0.26}$. It is observed that the error in $H_{0}$ has decreased. In Figure 4, $\alpha$CDM is seen to fit well with the observational data from Pantheon and is close to the $\Lambda$CDM model. Next, we included an additional 6 BAO data points and obtained the fitting results: $H_{0}=67.5^{+1.3}_{-1.6}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.2764\pm{0.0094}$, and $\alpha=0.33\pm{0.22}$. In Table 5, we can find with the increase in the observational data size, the constraints on the parameters $H_{0}$, $\Omega_{\rm{m0}}$, and $\alpha$ become progressively more accurate. To validate the model’s efficacy, we compared error bar plots for Hubble and Pantheon datasets with the $\Lambda$CDM model. Consistently, the model fits the observed dataset very well.
@@
From Table V, we observe that the fitted values of $H_{0}$ using different datasets are slight variations, but are relatively close, indicating that $H_{0}$ is not highly sensitive to the choice of dataset. The fitted values of $\Omega_{\rm{m}0}$ show minimal variation, ranging from 0.264 to 0.2764, suggesting low sensitivity to dataset choice. However, the fitted values of $\alpha$ vary more significantly across different datasets, ranging from 0.71 to 0.33, indicating higher sensitivity and greater uncertainty for this parameter. Compared to $\Lambda$CDM constrained from Pantheon+SH0ES Brout et al. (2022) and the model proposed in Li and Shafieloo (2019) with two parameters constrained from Pantheon+BAO data, the values of $H_{0}$ differ significantly from the values obtained here, meaning the results are sensitive to the choice of model parameters. Compared to $w$CDM constrained here and the model proposed in Myrzakulov et al. (2023) with three parameters constrained from Hubble+Pantheon+BAO data, the values of $H_{0}$ differ slightly from the values obtained here, implying the results may be not sensitive to the choice of model parameters.
@@ V Conclusions
In conclusion, we proposed a new Hubble parameterization method and constrained the model parameters using observations from Hubble, Pantheon, and BAO. This model is validated with Hubble, Pantheon, and BAO data, providing best-fit values for $H_{0}=67.5^{+1.3}_{-1.6}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.2764\pm{0.0094}$, and $\alpha=0.33\pm{0.22}$, consistent well with the Planck 2018 data ($H_{0}=67.4\pm 0.5$ Km/s/Mpc Aghanim et al. (2020)), but deviating from Cepheid-supernova observation ($H_{0}=73.04\pm 1.04$ $\text{km s}^{-1}\text{Mpc}^{-1}$ Riess et al. (2022)) by more than 4 $\sigma$. The Hubble tension was mainly triggered with the higher Hubble constant ($H_{0}=74.0\pm 1.04$ $\text{km s}^{-1}\text{Mpc}^{-1}$ Riess et al. (2022)) estimated from the local Cepheid–type SNe Ia distance ladder being at odds with the lower value extrapolated from CMB data, assuming the standard $\Lambda$CDM cosmological model ($H_{0}=67.4\pm 0.5$ $\text{km s}^{-1}\text{Mpc}^{-1}$ Aghanim et al. (2020)). However, our analyses indicate that the Hubble tension may also exist between the Hubble+Pantheon+BAO data and the measurements from local Cepheid–type Ia supernova distance ladder if taking the parametrization (6). We also constrained $w$CDM with the same datasets for comparison and found that $\alpha$CDM is consistent with $w$CDM within 1 $\sigma$ error, indicating that the impact of radiation is relatively weak.
Any potential biases or systematic errors in the observational data may have some impact on the results here. Completely eliminating biases and systematic errors in observational data is challenging and impossible. The datasets we used have undergone extensive cross-calibration and systematic error corrections by other researchers to minimize residual biases and observational errors Brout et al. (2022); Riess et al. (2022). Additionally, we accounted for these residual errors in our simulations and analyses by incorporating error terms in the Monte Carlo likelihood function and performing error propagation analysis. Using multiple datasets also aids in cross-verifying results, thereby reducing the influence of systematic errors from any single dataset. Table 5 provides the upper and lower limits for each parameter, demonstrating the sensitivity of our results to these errors. These bounds indicate that the impact of these errors is limited and does not significantly affect the overall conclusions of our analysis.